OpenStudy (anonymous):

Solve for x. e^4x=(1/64) so first I can do ln e ^4x=ln(1/64) right? I am then having trouble from here. Please help.

OpenStudy (campbell_st):

you need to take the base e log of both sides so you have \[\ln(e^{4x}) = \ln(\frac{1}{64})\]

OpenStudy (campbell_st):

the left hand side of the equation becomes 4x so you are solving \[4x = \ln(\frac{1}{64})\] you need to use the lag law for division and the law for ln(1) to get the answer

OpenStudy (anonymous):

When you say log law for division are you talking about log base b(x)-log base b (y)?

OpenStudy (campbell_st):

on the right hand side you are dividing 1 by 64 so you may choose to use 4x = ln(1) - ln(64) and then work from there

OpenStudy (campbell_st):

but you need to know ln(1)

OpenStudy (anonymous):

Ok so ln base e (1)=1 right? So I did 4x-1=ln base e 64

OpenStudy (anonymous):

ok wait so ln(1) is not the same as ln base e(1) correct?

OpenStudy (anonymous):

ln(1)=0 ?

OpenStudy (campbell_st):

not quite ln(1) = 0 so you would have 4x = 0 - ln(64) solve for x

OpenStudy (anonymous):

Use the property \[\log_{a}(a)^p = p \] ,,,, that is , when base and the final number are same, take power as the value..

OpenStudy (anonymous):

But I don't see the base and final number being the same. Right now I am at x=-ln(64)/4

OpenStudy (campbell_st):

this is what you need to solve \[4x = - \ln(64)\] just solve for x

OpenStudy (anonymous):

I meant that \[\log_{e} (e)^(4x) = 4x\] And then, you can directly write \[\log_{e} (1/64) = \ln (4)^(-3)\]